# Multimodal Trajectory Representation Learning for Travel Time Estimation

### v1

@@ Abstract.
Accurate travel time estimation (TTE) plays a crucial role in intelligent transportation systems. However, it remains challenging due to heterogeneous data sources and complex traffic dynamics. Moreover, conventional approaches typically convert trajectories into fixed-length representations, neglecting the inherent variability of real-world trajectories, which often leads to information loss or feature redundancy. To address these challenges, this paper introduces the Multimodal Dynamic Trajectory Integration (MDTI) framework–a novel multimodal trajectory representation learning approach that integrates GPS sequences, grid trajectories, and road network constraints to enhance TTE accuracy. MDTI employs modality-specific encoders and a cross-modal interaction module to capture complementary spatial, temporal, and topological semantics, while a dynamic trajectory modeling mechanism adaptively regulates information density for trajectories of varying lengths. Two self-supervised pretraining objectives, named contrastive alignment and masked language modeling, further strengthen multimodal consistency and contextual understanding. Extensive experiments on three real-world datasets demonstrate that MDTI consistently outperforms state-of-the-art baselines, confirming its robustness and strong generalization abilities. The code is publicly available at: https://github.com/freshhxy/MDTI/
Relevance Statement: This paper proposes a method to integrate open-source GPS sensor data, grid trajectories, and road network semantics for dynamic multimodal trajectory representation learning, addressing semantic fusion and representation challenges in large-scale mobility data on the web.
Trajectory representation learning, Multimodal data mining
journalyear: 2026isbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
With the rapid development of connected vehicle technologies, vast amounts of trajectory-based sensor data are continuously generated and shared on the Web. Within this context, travel time estimation (TTE) (Han et al., 2023) has emerged as a core task in traffic flow analysis. By transforming heterogeneous trajectory data into a unified indicator, TTE provides essential input for downstream decision-making processes such as route planning, traffic guidance, and system-wide control. TTE focuses on predicting the time required to traverse a specific route from an origin to a destination, while accounting for dynamic factors such as traffic conditions, road characteristics, and environmental influences (Liu et al., 2023). Accurate predictions enable more efficient route planning, congestion mitigation, and resource management in urban transportation systems. Core aspects of TTE include its dependence on real-time data, integration of multimodal traffic information, and adaptability to rapidly changing traffic patterns. The main challenge, however, lies in effectively modeling the inherent complexity of multimodal spatiotemporal traffic systems. This underscores the need for advanced approaches to improve the precision and robustness of TTE.
@@
Figure 1. Toy examples of trajectory data.
Trajectory representation learning (Li et al., 2018b; Yao et al., 2017b) provides an effective framework for modeling multimodal spatiotemporal data, playing a crucial role in reconstructing complex trajectory patterns and supporting TTE tasks. Early trajectory representation learning (TRL) approaches mainly employed univariate sequence-to-sequence architectures (Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017) to encode raw GPS trajectories (Fu and Lee, 2020; Li et al., 2018b; Yao et al., 2017b). In these methods, trajectories are modeled as sequences of time-stamped data points, where each point corresponds to a single variable such as latitude, longitude, or speed. Although such models effectively capture temporal dependencies and map sequential data into latent embeddings, they fall short in representing multimodal trajectory information, thereby limiting their applicability and robustness in real-world traffic scenarios.
@@
Another challenge in trajectory representation lies in converting dynamically varying trajectory samples into fixed-length feature vectors. For instance, GPS data collected from different vehicles often differ in spatiotemporal sampling frequencies (Andersen and Torp, 2017), sequence lengths (Wang et al., 2018a), and other factors. However, many existing methods address trajectory variability by setting a maximum sequence length (Jiang et al., 2023), padding shorter trajectories with zeros, and clipping longer ones for representation learning. However, these strategies often introduce redundant noise or cause information loss. As shown in Figure 1(b), when trajectories exceed the maximum threshold (top left), only the last segment is retained, preventing a full representation of the overall trajectory. Conversely, shorter trajectories (top right) are padded with redundant GPS points, which overemphasize local features such as the starting point and hinder effective representation learning. To overcome these issues, dynamic modeling techniques (bottom) can adaptively generate embeddings with differentiated information density and reduced redundancy as trajectory lengths vary. Therefore, the second challenge addressed in this work is to incorporate dynamic modeling into trajectory representation learning, ensuring fixed-length embeddings that better capture essential spatiotemporal patterns.
To address these challenges, we propose the Multimodal Dynamic Trajectory Integration Model (MDTI), which jointly models GPS sequences, grid trajectories, and road network constraints to comprehensively represent trajectory features for travel time estimation. The proposed framework incorporates a cross-modal interaction module that captures relationships among different modalities, enabling effective alignment and consistent semantic extraction. In addition, dedicated modality-specific encoders map heterogeneous inputs into a unified feature space, facilitating seamless multimodal fusion. Furthermore, a dynamic trajectory extraction module adaptively adjusts the information density of trajectory feature vectors according to trajectory length, preventing feature over-concentration in local embeddings and promoting balanced and discriminative representations. The main contributions of this work are summarized as follows:
@@
  • We design a cross-modal interaction module and a dynamic trajectory modeling module: the former aligns heterogeneous features and extracts consistent semantics across modalities, while the latter adaptively encodes trajectories of varying lengths, mitigating the information loss and redundancy inherent in fixed-length modeling.
@@
2. Related Work
2.1. Multimodal Trajectory Representation Learning
TRL encodes sequences of states, actions, or spatiotemporal points into compact vectors that preserve temporal structures and semantic information. Early work introduced Traj2vec (Yao et al., 2017b), which directly modeled raw GPS trajectories as point sequences to derive compact representations for downstream tasks. However, due to inherent issues in GPS data—such as positioning drift, U-turns, and redundant sampling—these approaches often produced unstable and poorly transferable representations.
To overcome these limitations, grid-based trajectory modeling methods have been proposed (Chang et al., 2023; Li et al., 2018c; Yao et al., 2022). By discretizing the spatial domain into grid cells and constructing graph-based structures, these approaches enhance the capacity for semantic and structural trajectory modeling. For instance, t2vec (Li et al., 2018c) employs an encoder–decoder architecture (Sutskever et al., 2014) to learn robust trajectory representations from noisy or low-quality data. E2DTC (Fang et al., 2021) introduces a self-training mechanism to capture latent spatial dependencies within grid-based embeddings. More recently, TrajCL (Chang et al., 2023) adopts contrastive learning to jointly model spatial correlations and structural relationships among trajectories.
In parallel, road-based trajectory modeling has emerged as another major direction. Using map-matching techniques (Lou et al., 2009; Yang and Gidofalvi, 2018), trajectories are aligned with road networks, which introduces topological constraints and enables richer representations of travel patterns. JCLRNT (Mao et al., 2022) generates contrastive samples between roads by leveraging contextual neighbors and constructing trajectory-level contrasts through detour path replacement. Start (Jiang et al., 2023) incorporates temporal regularities and travel semantics into a self-supervised framework, improving both representation quality and generalization. JGRM (Ma et al., 2024) further strengthens spatiotemporal road features by mapping raw GPS points to corresponding road segments. Building on this trend, GREEN (Zhou et al., 2025) integrates grid-based and road-based representations, exploiting their complementarity to yield more comprehensive trajectory embeddings and improved downstream task performance.Building on these advances, our study takes a further step by integrating GPS-based, grid-based, and road-based information into a unified multimodal framework, thereby addressing the limitations of single-modality representations for TTE.
2.2. Travel Time Estimation
TTE refers to predicting the time required to traverse a route or road segment from origin to destination. Existing studies can be broadly categorized into path-based methods and origin–destination (OD)-based methods.
Path-based methods use complete trajectories or trajectory segments as input, enabling the modeling of vehicle dynamics within the road network. These approaches typically leverage road segment features, trajectory sequences, or path structures to capture spatiotemporal dependencies. For instance, DeepTTE (Wang et al., 2018b) integrates graph convolutional networks with LSTM (Yao et al., 2017a) to model the spatiotemporal dependencies of GPS sequences. HierETA (Chen et al., 2022) introduces a multi-view hierarchical choice model for fine-grained and interpretable TTE. MulT-TTE (Liao et al., 2024) jointly models trajectory, attribute, and semantic sequences, and further applies self-supervised learning to improve generalization, yielding significant gains in prediction accuracy.
OD-based methods, by contrast, rely only on origin–destination pairs and departure times, avoiding the need for full trajectory inputs. TEMP (Wang et al., 2019) estimates travel times using only OD pairs and departure information. Extending this, MURAT (Li et al., 2018a) incorporates road network structures and driver behavior to enrich spatiotemporal embeddings. More recently, DutyTTE (Mao et al., 2025) combines reinforcement learning with a mixture-of-experts framework to optimize path prediction and capture segment-level uncertainty, enabling confidence interval estimation.
Building on these directions, our work is more closely aligned with path-based approaches, as it exploits the spatiotemporal information embedded in complete trajectories to achieve more accurate and robust TTE.
3. Overview
3.1. Definitions
@@
where $x_{i}$ and $y_{i}$ denote longitude and latitude, and $t_{i}$ denotes the timestamp. Here, $T$ is the total number of points in the trajectory.
Definition 2 (Grid Trajectory). A grid trajectory $\mathcal{T}^{grid}$ maps GPS points onto a discretized spatial grid. Let the region of interest be partitioned into $M\times N$ cells, each representing a spatial unit. The trajectory is then expressed as a sequence of grid cell identifiers:
$\mathcal{T}^{grid}=\{g_{t}\}_{t=1}^{T},\quad g_{t}\in\{1,\ldots,M\times N\},$
where $g_{t}$ denotes the identifier (ID) of the grid cell visited at time $t$.
Definition 3 (Road Network). A road network is represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{|\mathcal{V}|}\}$ denotes the set of vertices, and each vertex $v_{i}\in\mathcal{V}$ corresponds to a distinct road segment. $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of directed edges, where each edge $e_{ij}=(v_{i},v_{j})\in\mathcal{E}$ represents the navigable connection from road segment $v_{i}$ to $v_{j}$.
@@
$\mathcal{T}^{road}=\{v_{t}\}_{t=1}^{T},\quad v_{t}\in\mathcal{V},$
@@ 4.1. Grid Encoder
(5) $H^{(2)}=\sigma\!\left(\text{GATConv}^{(2)}\!\left(\sigma\!\left(\text{GATConv}^{(1)}(X,A)\right),A\right)\right),$
and we define $H=H^{(2)}$ as the final grid-level representation. Before applying the projection, we aggregate the node embeddings of all grid cells into the matrix form $H$, which serves as the grid-level embedding map for subsequent temporal modeling. The embedding sequence is then projected as
(6) $G_{\text{emb}}=\text{Proj}(H)\in\mathbb{R}^{T\times D},$
where $\text{Proj}(\cdot)$ denotes a linear projection, $T$ is the number of temporal segments, and $D$ is the embedding dimension of the Transformer input. The resulting representation $G_{\text{emb}}=\{G_{0},G_{1},\dots,G_{T}\}$ serves as the spatio-temporal embedding sequence of grid nodes, with $G_{0}$ corresponding to the [CLS] token, and the remaining tokens representing different time segments of the trajectory are mapped onto the spatial grids. This sequence embedding is then used as input for subsequent trajectory sequence modeling and Transformer-based encoding.
@@ 4.2. GPS Encoder
After obtaining the structured motion information, we construct a natural-language prompt that integrates the temporal indices, cumulative trend, and pattern similarity results from Equation 8. The prompt for the $i$-th trajectory segment is formulated as:
@@
By aggregating all entities in a batch, we obtain the GPS modality representation:
(11) $Z=\{z_{1},z_{2},\ldots,z_{N}\},$
which captures the real motion patterns of the vehicle in continuous space.
4.3. Dynamic FusionAlign
Traditional cross-attention fusion mechanisms enable interaction between modalities but suffer from inherent asymmetry: one modality dominates as the query source while the other serves as the key-value counterpart, leading to imbalanced fusion and difficulty adapting to varying sequence lengths and structures. To address this, we propose a dynamic fusion mechanism that replaces the dominant–subordinate paradigm with symmetric and adaptive deep fusion, allowing bidirectional feature enhancement through fine-grained alignment and maximizing complementary information integration. In our framework, the trajectory representation generated by the GPS encoder acts as a prompt that guides and reinforces sequence learning in the grid encoder. Unlike conventional methods that simply concatenate or append GPS features to downstream predictors, our Dynamic Fusion module operates within the encoding stage, establishing early cross-modal correspondences between grid-based and GPS-based representations to ensure temporal and spatial consistency across modalities.
@@ 4.3. Dynamic FusionAlign
Finally, the aligned GPS sequence $e^{\prime}$ is injected into the grid sequence via a residual fusion mechanism:
@@ 4.4. Road Encoder
Unlike the grid encoder, where the neighborhood $\mathcal{N}(i)$ is defined by fixed spatial proximity in a $3\times 3$ grid, the road encoder constructs $\mathcal{N}(i)$ according to the topological connectivity of the road network. This allows message passing along physically connected road segments rather than adjacent spatial grids. Through this topology-aware aggregation, each node representation integrates its own attributes and the structural dependencies of neighboring roads, producing road-aware graph embeddings $g_{\text{emb}}$.
On the sequence side, we construct the road sequence embedding table by augmenting $\mathbf{R}_{\text{emb}}$ with three special tokens [PAD], [CLS], and [MASK]. Given a trajectory $\mathbf{R}_{\text{traj}}$, we obtain the corresponding sequence embeddings and further add temporal embeddings (week index and minute index), type embeddings, and positional encodings $\mathrm{PE}(\cdot)$ to preserve ordering. A padding mask ensures that attention is only applied to valid road segments.
@@ 4.5. Multimodal Interactor
Specifically, given a road sequence $X\in\mathbb{R}^{B\times L_{x}\times d}$ and a dynamically fused grid representation $M\in\mathbb{R}^{B\times L_{y}\times d}$, we compute the cross-modal attention as:
(16) $\text{Attn}(X,M)=\text{Softmax}\!\left(\frac{(XW_{Q})(MW_{K})^{\top}}{\sqrt{d_{k}}}+M_{s}\right)(MW_{V}),$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{k}}$ are learnable projection matrices, and $M_{s}$ denotes the spatial alignment bias. In this design, the road sequence $X$ acts as the query, actively retrieving relevant contextual information, while the fused grid representation $M$ serves as the key and value, providing rich spatial semantics. This formulation enables adaptive and context-aware feature interaction between modalities, guided by the trajectory structure.
The fused output is then refined through a residual connection followed by a feed-forward network (FFN):
(17) $\hat{X}=\text{FFN}\big(X+\text{Attn}(X,M)\big),$

### v2

@@ Abstract.
Accurate travel time estimation (TTE) plays a crucial role in intelligent transportation systems. However, it remains challenging due to heterogeneous data sources and complex traffic dynamics. Moreover, traditional approaches typically convert trajectory data into fixed-length representations. This overlooks the inherent variability of real-world motion patterns, often resulting in information loss and redundancy. To address these challenges, this paper introduces the Multimodal Dynamic Trajectory Integration (MDTI) framework–a novel multimodal trajectory representation learning approach that integrates GPS sequences, grid trajectories, and road network constraints to enhance the performance of TTE. MDTI employs modality-specific encoders and a multimodal fusion module to capture complementary spatial, temporal, and topological semantics, while a dynamic trajectory modeling mechanism adaptively regulates information density for trajectories of varying lengths. Two self-supervised pretraining objectives, named contrastive alignment and masked language modeling, further strengthen multimodal consistency and contextual understanding. Extensive experiments on three real-world datasets demonstrate that MDTI consistently outperforms state-of-the-art baselines, confirming its robustness and strong generalization abilities. The code is publicly available at: https://github.com/City-Computing/MDTI.
Trajectory representation learning, Multimodal data mining, Travel time estimation
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792388isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Spatial-temporal systemsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
With the rapid development of connected vehicle technologies, vast amounts of trajectory-based sensor data are continuously generated and shared on the Web. Within this context, travel time estimation (TTE) (Han et al., 2023a) has emerged as a core task in traffic flow analysis (Zhang et al., 2017; Zheng, 2015; Liu et al., 2024; Han et al., 2024). By transforming heterogeneous trajectory data into a unified indicator, TTE provides essential input for downstream decision-making processes such as route planning, traffic guidance, and can also support signal control (Han et al., 2023b). TTE focuses on predicting the time required to traverse a specific route from an origin to a destination, while accounting for dynamic factors such as traffic conditions (Shen et al., 2021), road characteristics, and environmental influences (Liu et al., 2023). Accurate estimation enables more efficient route planning (Chen et al., 2025), congestion mitigation (Han et al., 2020), and signal control (Han et al., 2023b) in urban transportation systems. Core aspects of TTE include its dependence on real-time data, integration of multimodal traffic information, and adaptability to rapidly changing traffic patterns. The main challenge, however, lies in effectively modeling the inherent complexity of multimodal spatiotemporal traffic systems. This underscores the need for advanced approaches to improve the precision and robustness of TTE.
@@
Figure 1. Toy examples of online trajectory data.
Trajectory representation learning (Li et al., 2018; Yao et al., 2017a) provides an effective framework for modeling multimodal spatiotemporal data, playing a crucial role in reconstructing complex trajectory patterns and supporting TTE tasks. Early trajectory representation learning (TRL) approaches mainly employed univariate sequence-to-sequence architectures (Hochreiter and Schmidhuber, 1997; Vaswani et al., 2017) to encode raw GPS trajectories (Fu and Lee, 2020; Han et al., 2025). In these methods, trajectories are modeled as sequences of time-stamped data points, where each point contains a spatial coordinate pair (latitude and longitude), optionally accompanied by other attributes such as speed. Although such models effectively capture temporal dependencies and map sequential data into latent embeddings, they fall short in representing multimodal trajectory information, thereby limiting their applicability and robustness in real-world traffic scenarios.
@@
Another challenge in trajectory representation lies in converting dynamically varying trajectory samples into fixed-length feature vectors. For instance, GPS data collected from different vehicles often differ in spatiotemporal sampling frequencies (Andersen and Torp, 2017), sequence lengths (Wang et al., 2018b), and other factors. Many existing methods address trajectory variability by setting a maximum sequence length (Jiang et al., 2023), padding shorter trajectories with zeros, and clipping longer ones for representation learning. However, these strategies often introduce redundant noise (Chen et al., 2024) or cause information loss. As shown in Figure 1(b), when trajectories exceed the maximum threshold (top left), only the last segment is retained, preventing a full representation of the overall trajectory. Conversely, shorter trajectories (top right) are padded with redundant GPS points, which overemphasize local features such as the starting point and hinder effective representation learning. To overcome these issues, dynamic modeling techniques (bottom) can adaptively generate embeddings with differentiated information density and reduced redundancy as trajectory lengths (Berndt and Clifford, 1994; Chen and Ng, 2004; Chen et al., 2005) vary. Therefore, The second challenge is encoding trajectories of varying lengths and structures into uniform fixed-length feature representations.
To address these challenges, we propose the Multimodal Dynamic Trajectory Integration Model (MDTI), which jointly models GPS sequences, grid trajectories, and road network constraints to comprehensively represent trajectory features for travel time estimation. The proposed framework incorporates a multimodal fusion module that captures relationships among different modalities, enabling effective alignment and consistent semantic extraction. In addition, dedicated modality-specific encoders map heterogeneous inputs into a unified feature space, facilitating seamless multimodal fusion. Furthermore, a dynamic trajectory extraction module adaptively adjusts the information density of trajectory feature vectors according to trajectory length, preventing feature over-concentration in local embeddings and promoting balanced and discriminative representations. The main contributions of this work are summarized as follows:
@@
  • We designed a multimodal fusion module that not only aligns heterogeneous features and extracts consistent cross-modal semantics but also adaptively encodes trajectories of varying lengths, mitigating the inherent information loss and redundancy issues in fixed-length modeling.
@@
2. Preliminary
@@
where $x_{i}$ and $y_{i}$ denote longitude and latitude, and $t_{i}$ denotes the timestamp. Here, $T$ is the total number of points in the trajectory. Definition 2 (Grid Trajectory). A grid trajectory $\mathcal{T}^{grid}$ maps GPS points onto a discretized spatial grid. Let the region of interest be partitioned into $M\times N$ cells, each representing a spatial unit. The trajectory is then expressed as a sequence of grid cell identifiers:
(2) $\mathcal{T}^{grid}=\{g_{t}\}_{t=1}^{T},\quad g_{t}\in\{1,\ldots,M\times N\},$
where $g_{t}$ denotes the identifier (ID) of the grid cell visited at time $t$. Definition 3 (Road Network). A road network is represented as a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},v_{2},\ldots,v_{|\mathcal{V}|}\}$ denotes the set of vertices, and each vertex $v_{i}\in\mathcal{V}$ corresponds to a distinct road segment. $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of directed edges, where each edge $e_{ij}=(v_{i},v_{j})\in\mathcal{E}$ represents the navigable connection from road segment $v_{i}$ to $v_{j}$.
@@
(3) $\mathcal{T}^{road}=\{v_{t}\}_{t=1}^{T},\quad v_{t}\in\mathcal{V},$
@@ 3.1. Grid Encoder
(7) $H^{(2)}=\sigma\!\left(\text{GATConv}^{(2)}\!\left(\sigma\!\left(\text{GATConv}^{(1)}(X,A)\right),A\right)\right).$
Before applying the projection, we aggregate the node embeddings of all grid cells into the matrix form $H$, which serves as the grid-level embedding map for subsequent temporal modeling. The embedding sequence is then projected as
(8) $G=\text{Proj}(H)\in\mathbb{R}^{T\times D},$
where $\text{Proj}(\cdot)$ denotes a linear projection, $T$ is the number of temporal segments, and $D$ is the Transformer input embedding dimension.
@@ 3.2. GPS Encoder
After obtaining the structured motion information, we construct a natural-language prompt that integrates the temporal indices, cumulative trend, and pattern similarity results from Eq. 10. The prompt for the $i$-th trajectory segment is formulated as:
@@
This design combines prototype-based alignment, structured prompt generation, and representation learning with large-scale language models. It enables GPS trajectories to provide fine-grained motion semantics while leveraging the model’s capacity to capture semantic dependencies and extract latent relationships across trajectories, thereby enhancing the overall framework’s expressive power.
3.3. Dynamic Fusion
Traditional cross-attention fusion mechanisms enable interaction between modalities but suffer from inherent asymmetry: one modality dominates as the query source while the other serves as the key-value counterpart, leading to imbalanced fusion and difficulty adapting to varying sequence lengths and structures. To address this, we propose a dynamic fusion mechanism that replaces the dominant–subordinate paradigm with symmetric and adaptive deep fusion, allowing bidirectional feature enhancement through fine-grained alignment and maximizing complementary information integration. In our framework, the trajectory representation generated by the GPS encoder acts as a prompt that guides and reinforces sequence learning in the grid encoder. Unlike conventional methods that simply concatenate or append GPS features to downstream predictors, our Dynamic Fusion module operates within the encoding stage, establishing early cross-modal correspondences between grid-based and GPS-based representations to ensure spatial and temporal consistency across modalities.
@@ 3.3. Dynamic Fusion
Finally, the aligned GPS sequence $z^{\prime}$ is injected into the grid sequence via a residual fusion mechanism:
@@ 3.4. Road Encoder
Unlike the grid encoder, where the neighborhood $\mathcal{N}(i)$ is defined by fixed spatial proximity in a $3\times 3$ grid, the road encoder constructs $\mathcal{N}(i)$ according to the topological connectivity of the road network. This allows message passing along physically connected road segments rather than adjacent spatial grids. Through this topology-aware aggregation, each node representation integrates its own attributes and the structural dependencies of neighboring roads, producing road-aware graph embeddings.
On the sequence side, we construct the road sequence embedding table by augmenting three special tokens [PAD], [CLS], and [MASK]. Given a trajectory, we obtain the corresponding sequence embeddings and further add temporal embeddings (week index and minute index), type embeddings, and positional encodings $\mathrm{PE}(\cdot)$ to preserve ordering. A padding mask ensures that attention is only applied to valid road segments.
@@ 3.5. Multimodal Interactor
Specifically, given a fused grid representation $\tilde{G}\in\mathbb{R}^{B\times L_{g}\times d}$ and a road sequence $\mathbf{R}\in\mathbb{R}^{B\times L_{r}\times d}$, we compute the cross-modal attention as:
(17) $\text{Attn}(\tilde{G},\mathbf{R})=\text{Softmax}\left(\frac{(\tilde{G}W_{Q})(\mathbf{R}W_{K})^{\top}}{\sqrt{d_{k}}}+M_{s}\right)(\mathbf{R}W_{V}),$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{k}}$ are learnable projection matrices, and $M_{s}$ denotes the spatial alignment bias. In this design, the road sequence $\mathbf{R}$ acts as the query, actively retrieving relevant contextual information, while the fused grid representation $\tilde{G}$ serves as the key and value, providing rich spatial semantics. This formulation enables adaptive and context-aware feature interaction between modalities, guided by the trajectory structure.
The fused output is then refined through a residual connection followed by a feed-forward network:
(18) $\hat{X}=\text{FFN}\left(\tilde{G}+\text{Attn}(\tilde{G},\mathbf{R})\right),$